# Solving General-Utility Markov Decision Processes in the Single-Trial Regime with Online Planning

### v1

@@ 3.1 Non-Markovianity matters
The result above shows that, in general, the class of stationary policies is strictly dominated by the class of non-stationary policies, which is, in turn, strictly dominated by the class of non-Markovian policies. Hence, non-Markovianity matters and we must focus our attention to history-dependent policies when optimizing GUMDPs in the single-trial regime. Our Theo. 1 extends the result in Mutti et al. (2023), which considers finite-horizon GUMDPs, to the infinite-horizon discounted setting.
@@ 3.2 Computing (approximately) optimal policies by resorting to $F_{1,H}$
The result below (proof in Appendix B.2) establishes that the regret $\mathcal{R}(\pi)$ of any policy $\pi\in\Pi_{\text{NM}}$ can be upper bounded, up to a constant, by the regret of policy $\pi$ for the single-trial truncated objective.
Proposition 1 (Regret decomposition).
@@
$\mathcal{R}(\pi)\leq\underbrace{F_{1,H}(\pi)-\min_{\pi_{H}\in\Pi_{\textnormal{NM}}}\left\{F_{1,H}(\pi_{H})\right\}}_{\text{$=\mathcal{R}_{H}(\pi)$}}+8L\gamma^{H},$ (7)
where $\mathcal{R}_{H}(\pi)$ is the regret of policy $\pi$ under the single-trial truncated objective with horizon $H$.
Intuitively, the proposition above shows that we can resort to the single-trial truncated objective, as introduced in (6), to find an approximately optimal policy for the original objective defined in (5), up to any desired tolerance. In particular, if $\pi$ is the optimal policy for the single-trial truncated objective, i.e., $\mathcal{R}_{H}(\pi)=0$, then it holds that $\mathcal{R}(\pi)\leq 8L\gamma^{H}$, which can be made arbitrarily small by tuning our truncation horizon $H$. Therefore, we focus our attention on how to compute
@@
in order to keep the truncated regret term $\mathcal{R}_{H}(\pi)$ low, which we investigate in the next section.
@@ 3.3 The occupancy MDP: Casting $\mathcal{M}_{f}$ as a standard MDP
To derive our planning algorithms for solving GUMDPs in the single-trial truncated setting, we derive a finite-horizon MDP based on the original GUMDP. In particular, we consider the occupancy MDP defined by the tuple $\mathcal{M}_{\text{O}}=\{\mathcal{S}_{\text{O}},\mathcal{A}_{\text{O}},\{{\boldsymbol{P}}_{\text{O}}^{a}\},{\boldsymbol{p}}_{0,\text{O}},c_{\text{O}},H\}$, where $\mathcal{S}_{\text{O}}=\mathcal{S}\times\mathcal{O}$ is the discrete state space and
$\mathcal{O}$ =$=\Bigg{\{}{\boldsymbol{o}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}:o(s,a)=\sum_{t=0}^{l-1}\gamma^{t}\mathbf{1}(s_{t}=s,a_{t}=a),\forall s\in\mathcal{S},a\in\mathcal{A},$
@@
We denote a state of the occupancy MDP with the tuple $\{s,{\boldsymbol{o}}\}$, where $s\in\mathcal{S}$ is a state from the original GUMDP and ${\boldsymbol{o}}\in\mathcal{O}$ is a $|\mathcal{S}||\mathcal{A}|$-dimensional vector that keeps track of the running occupancy of the agent up to a given timestep. Intuitively, the running occupancy records the empirical occupancy, as defined in (3), observed by the agent up to any timestep. We let $\mathcal{A}_{\text{O}}=\mathcal{A}$ be the action space. We define ${\boldsymbol{p}}_{0,\text{O}}$ such that $p_{0,\text{O}}(\{s,{\boldsymbol{o}}\})=p_{0}(s)$ if ${\boldsymbol{o}}=[0,\ldots,0]$ and zero otherwise. The dynamics are as follows: (i) component $\mathrm{s}_{t+1}\sim P^{\mathrm{a}_{t}}(\cdot|\mathrm{s}_{t})$ evolves according to the dynamics of the original GUMDP; and (ii) the running occupancy evolves deterministically as $o_{t+1}(s,a)=\gamma^{t}+o_{t}(s,a)$ if $s=\mathrm{s}_{t}$ and $a=\mathrm{a}_{t}$, and $o_{t+1}(s,a)=o_{t}(s,a)$ otherwise. We emphasize that we do not need to incorporate the timestep in the state of the occupancy MDP since it can be inferred from the running occupancy by summing its entries. Finally, $H\in\mathbb{N}$ denotes the horizon of the MDP and the cost function $c_{\text{O}}:\mathcal{S}\times\mathcal{O}\rightarrow\mathbb{R}$ is defined as
$c_{\text{O}}(\{s,{\boldsymbol{o}}\})=\begin{cases}0&\text{if $t<H$},\\ f\left(\frac{1-\gamma}{1-\gamma^{H}}{\boldsymbol{o}}\right)&\text{if $t=H$}.\end{cases}$
@@
where the expectation above is taken with respect to the random sequence of states ($(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\ldots,\{\mathrm{s}_{H},\mathbf{o}_{H}\})$ under policy $\pi_{\text{O}}$. We let $J_{\text{O}}^{*}=\min_{\pi_{\text{O}}\in\Pi_{\text{S}}^{\text{D}}}J_{\text{O}}(\pi_{\text{O}})$ be the optimal cumulative cost for $\mathcal{M}_{\text{O}}$. We also note that the occupancy MDP possesses well-defined (optimal) value and action-value functions, which can be shown to satisfy standard Bellman equations (Appendix B.3).
@@
There exists a one-to-one mapping between histories $h_{l}=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{l})\in\mathcal{S}\times(\mathcal{S}\times\mathcal{A})^{l}$ in $\mathcal{M}_{f}$, with $10\leq l\leq H-$ 1, and states $\{s,{\boldsymbol{o}}\}\in\mathcal{S}\times\mathcal{O}$ in $\mathcal{M}_{\text{O}}$.
@@
The problem of finding a policy $\pi\in\Pi_{\textnormal{NM}}$ satisfying $\mathcal{R}_{H}(\pi)\leq\epsilon$, for any $\epsilon\in\mathbb{R}_{0}^{+}$, can be reduced to the problem of finding a policy $\pi_{\text{O}}\in\Pi_{\textnormal{S}}$ satisfying $J_{\text{O}}(\pi_{\text{O}})-J_{\text{O}}^{*}\leq\epsilon$. In particular, if $\pi^{*}_{\text{O}}=\operatorname*{arg\,min}_{\pi_{\text{O}}\in\Pi_{\textnormal{S}}}J_{\text{O}}(\pi_{\text{O}})$, then the corresponding non-Markovian policy $\pi$ in $\mathcal{M}_{f}$ satisfies $\mathcal{R}_{H}(\pi)=0$. Finally, it holds that $\mathcal{R}_{H}(\pi)=J_{\text{O}}(\pi_{\text{O}})-J^{*}_{\text{O}}$, where $\pi_{O}$ is the stationary policy for $\mathcal{M}_{\text{O}}$ associated with the non-Markovian policy $\pi$ for $\mathcal{M}_{f}$.
Intuitively, the result above tells us that it suffices to search for an approximately stationary optimal policy for $\mathcal{M}_{\text{O}}$, since such a policy corresponds to a non-Markovian policy that is approximately optimal for $\mathcal{M}_{f}$. In particular, an approximately optimal policy for $\mathcal{M}_{\text{O}}$ can be seen as a non-Markovian policy for $\mathcal{M}_{f}$ that compresses the history up to any timestep into a running occupancy. This result demonstrates that maintaining the running occupancy up to any timestep is sufficient to achieve optimal behavior in the single-trial truncated-horizon regime (8).
@@
Since the class of policies $\Pi_{\textnormal{S}}^{\textnormal{D}}$ suffices for optimality in standard MDPs (Puterman, 2014), we know that at least one stationary deterministic policy $\pi_{\text{O}}\in\Pi_{\textnormal{S}}^{\textnormal{D}}$ for $\mathcal{M}_{\text{O}}$ satisfies $J_{\text{O}}(\pi_{\text{O}})-J_{\text{O}}^{*}=0$, i.e., $\pi_{\text{O}}$ is optimal for $\mathcal{M}_{\text{O}}$. In light of Theo. 2 and Prop. 2, this implies that the corresponding non-Markovian policy $\pi\in\Pi_{\textnormal{NM}}$ for $\mathcal{M}_{f}$, which is deterministic, satisfies $\mathcal{R}_{H}(\pi)=0$, i.e., $\pi$ is optimal for $\mathcal{M}_{f}$. Thus, we can focus our attention on deterministic non-Markovian policies when solving $\mathcal{M}_{f}$ with objective $F_{1,H}$, i.e., for any GUMDP and horizon $H\in\mathbb{N}$, it holds that $\min_{\pi\in\Pi_{\textnormal{NM}}}F_{1,H}(\pi)=\min_{\pi\in\Pi_{\textnormal{NM}}^{\textnormal{D}}}F_{1,H}(\pi).$
@@
3.4 The computational complexity of policy optimization in the single-trial regime
@@
(Complete proof in Appendix B.6) We reduce the subset sum problem to the policy existence problem in GUMDPs with objective $F_{1,H}$. The subset sum problem asks whether, given a set $\mathcal{N}=\{n_{0},n_{1},\ldots,n_{N-1}\}$ of $N$ non-negative integers and a target sum $k\in\mathbb{N}$, there exists a subset of $\mathcal{N}$ whose elements sum to $k$. We map every instance of the subset sum problem as a GUMDP such that at each state $s_{i}$, for $i\in\{0,\ldots,N-1\}$, the policy $\pi\in\Pi_{\text{NM}}^{\text{D}}$ needs to decide between selecting: (i) $a_{\text{include}}$, thereby including $n_{i}$ in the sum; or (ii) $a_{\text{not-include}}$, thereby excluding $n_{i}$ from the sum. Then, we set $H\geq N$ and let $f({\boldsymbol{d}})=({\boldsymbol{n}}^{\top}{\boldsymbol{d}}-k)^{2}$, where ${\boldsymbol{d}}$ denotes a discounted occupancy that captures information regarding the actions selected by the agent at each state $s_{i}$. We construct vector ${\boldsymbol{n}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ such that ${\boldsymbol{n}}^{\top}{\boldsymbol{d}}$ equals the sum of the numbers selected by the policy. With this construction, the objective satisfies $f({\boldsymbol{d}})=0$ if and only if the sum of the selected numbers equals $k$. By setting $\lambda=0$, we are asking whether there exists a policy such that $F_{1,H}(\pi)\leq 0$. Since $f({\boldsymbol{d}})=0$ if and only if the selected numbers sum to $k$, the reduction is complete. ∎
@@
With the above hardness result in mind, we next explore how to develop practical planning algorithms for our problem. Naturally, in a worst-case sense, these algorithms may require a non-polynomial number of steps to retrieve the optimal policy. Nevertheless, our results show it is possible to develop practical algorithms that display superior performance in comparison to relevant baselines.
@@ 4 Online Planning for GUMDPs in the Single-Trial Regime
As previously shown, solving a GUMDP in the single-trial setting is closely related to solving a corresponding occupancy MDP. This connection allows us to employ an online planning approach in which, at any timestep $t\in\{0,\ldots,H-1\}$ of the interaction with the occupancy MDP, the algorithm receives the current state $\{s_{t},{\boldsymbol{o}}_{t}\}$. The online planner then expands a look-ahead search tree where the root node corresponds to state $\{s_{t},{\boldsymbol{o}}_{t}\}$. After a given number of iterations, the planning algorithm selects an action to execute in the environment; depending on the selected action, the environment evolves to a new state, and the process repeats until timestep $H$. This online planning strategy is particularly effective, as it allows computational resources to be focused on computing (approximately) optimal actions only along the specific trajectory experienced by the agent. This avoids the prohibitive cost of computing an optimal policy for every state of the GUMDP, whose state space is typically very large.

### v2

@@ 3.1 Non-Markovianity matters
There exists a GUMDP $\mathcal{M}_{f}$ with $\gamma\in(0,1)$ and $L$-Lipschitz convex objective such that (lower is better):
@@
The result above shows that, in general, the class of stationary policies is strictly dominated by the class of non-stationary policies, which is, in turn, strictly dominated by the class of non-Markovian policies. Hence, non-Markovianity matters, and we must focus our attention on history-dependent policies. Our Theo. 1 extends the result in Mutti et al. (2023), which considers finite-horizon GUMDPs, to the infinite-horizon discounted setting.
@@ 3.2 Computing (approximately) optimal policies by resorting to $F_{1,H}$
The result below (proof in Appendix B.2) establishes that the optimality gap $\mathrm{OptGap}(\pi)$ of any policy $\pi\in\Pi_{\text{NM}}$ can be upper bounded, up to a constant, by the optimality gap of policy $\pi$ for the single-trial truncated objective.
Proposition 1 (Optimality gap decomposition).
@@
$\mathrm{OptGap}(\pi)\leq\underbrace{F_{1,H}(\pi)-\min_{\pi_{H}\in\Pi_{\textnormal{NM}}}\left\{F_{1,H}(\pi_{H})\right\}}_{\text{$=\mathrm{OptGap}_{H}(\pi)$}}+8L\gamma^{H},$ (7)
where $\mathrm{OptGap}_{H}(\pi)$ is the optimality gap of policy $\pi$ under the single-trial truncated objective with horizon $H$.
Intuitively, the proposition above shows that we can resort to the single-trial truncated objective, as introduced in equation 6, to find an approximately optimal policy for the original objective defined in equation 5, up to any desired tolerance. In particular, if $\pi$ is the optimal policy for the single-trial truncated objective, i.e., $\mathrm{OptGap}_{H}(\pi)=0$, then it holds that $\mathrm{OptGap}(\pi)\leq 8L\gamma^{H}$, which can be made arbitrarily small by tuning our truncation horizon $H$. Therefore, we focus our attention on finding
@@
in order to keep the truncated optimality gap term $\mathrm{OptGap}_{H}(\pi)$ low, which we investigate in the next section.
@@ 3.3 The occupancy MDP: Casting $\mathcal{M}_{f}$ as a standard MDP
To derive our planning algorithms for solving GUMDPs in the single-trial truncated setting, we derive a finite-horizon MDP based on the original GUMDP. In particular, we consider the occupancy MDP defined by the tuple $\mathcal{M}_{\text{O}}=\{\mathcal{S}_{\text{O}},\mathcal{A}_{\text{O}},\{{\bm{P}}_{\text{O}}^{a}\},{\bm{p}}_{0,\text{O}},c_{\text{O}},H\}$, where $\mathcal{S}_{\text{O}}=\mathcal{S}\times\mathcal{O}$ is the discrete state space and
$\mathcal{O}$ $=\Bigg\{{\bm{o}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}:o(s,a)=\sum_{t=0}^{l-1}\gamma^{t}\mathbf{1}(s_{t}=s,a_{t}=a),\forall s\in\mathcal{S},a\in\mathcal{A},$
@@
We denote a state of the occupancy MDP with the tuple $\{s,{\bm{o}}\}$, where $s\in\mathcal{S}$ is a state from the original GUMDP and ${\bm{o}}\in\mathcal{O}$ is a $|\mathcal{S}||\mathcal{A}|$-dimensional vector that keeps track of the running occupancy of the agent up to a given timestep. Intuitively, the running occupancy records the empirical occupancy, as defined in equation 3, observed by the agent up to any timestep. We let $\mathcal{A}_{\text{O}}=\mathcal{A}$ be the action space. We define ${\bm{p}}_{0,\text{O}}$ such that $p_{0,\text{O}}(\{s,{\bm{o}}\})=p_{0}(s)$ if ${\bm{o}}=[0,\ldots,0]$ and zero otherwise. The dynamics are as follows: (i) component $\mathrm{s}_{t+1}\sim P^{\mathrm{a}_{t}}(\cdot|\mathrm{s}_{t})$ evolves according to the dynamics of the original GUMDP; and (ii) the running occupancy evolves deterministically as $o_{t+1}(s,a)=\gamma^{t}+o_{t}(s,a)$ if $s=\mathrm{s}_{t}$ and $a=\mathrm{a}_{t}$, and $o_{t+1}(s,a)=o_{t}(s,a)$ otherwise. We emphasize that we do not need to incorporate the timestep in the state of the occupancy MDP since it can be inferred from the running occupancy by summing its entries. Finally, $H\in\mathbb{N}$ denotes the horizon of the MDP and the cost function $c_{\text{O}}:\mathcal{S}\times\mathcal{O}\rightarrow\mathbb{R}$ is defined as
$c_{\text{O}}(\{s,{\bm{o}}\})=\begin{cases}0&\text{if $t<H$},\\ f\left(\frac{1-\gamma}{1-\gamma^{H}}{\bm{o}}\right)&\text{if $t=H$}.\end{cases}$
@@
where the expectation above is taken with respect to the random sequence of states $(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\ldots,\{\mathrm{s}_{H},\mathbf{o}_{H}\})$ under policy $\pi_{\text{O}}$. We let $J_{\text{O}}^{*}=\min_{\pi_{\text{O}}\in\Pi_{\text{S}}^{\text{D}}}J_{\text{O}}(\pi_{\text{O}})$ be the optimal cumulative cost for $\mathcal{M}_{\text{O}}$. We also note that the occupancy MDP possesses well-defined (optimal) value and action-value functions, which can be shown to satisfy standard Bellman equations (Appendix B.3).
@@
There exists a one-to-one mapping between histories $h_{l}=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{l})\in\mathcal{S}\times(\mathcal{S}\times\mathcal{A})^{l}$ in $\mathcal{M}_{f}$, with $0\leq l\leq H-1$, and states $\{s,{\bm{o}}\}\in\mathcal{S}\times\mathcal{O}$ in $\mathcal{M}_{\text{O}}$.
@@
The problem of finding a policy $\pi\in\Pi_{\textnormal{NM}}$ satisfying $\mathrm{OptGap}_{H}(\pi)\leq\epsilon$, for any $\epsilon\in\mathbb{R}_{0}^{+}$, can be reduced to the problem of finding a policy $\pi_{\text{O}}\in\Pi_{\textnormal{S}}$ satisfying $J_{\text{O}}(\pi_{\text{O}})-J_{\text{O}}^{*}\leq\epsilon$. In particular, if $\pi^{*}_{\text{O}}=\operatorname*{arg\,min}_{\pi_{\text{O}}\in\Pi_{\textnormal{S}}}J_{\text{O}}(\pi_{\text{O}})$, then the corresponding non-Markovian policy $\pi$ in $\mathcal{M}_{f}$ satisfies $\mathrm{OptGap}_{H}(\pi)=0$. Finally, it holds that $\mathrm{OptGap}_{H}(\pi)=J_{\text{O}}(\pi_{\text{O}})-J^{*}_{\text{O}}$, where $\pi_{O}$ is the stationary policy for $\mathcal{M}_{\text{O}}$ associated with the non-Markovian policy $\pi$ for $\mathcal{M}_{f}$.
Intuitively, the result above tells us that it suffices to search for an approximately stationary optimal policy for $\mathcal{M}_{\text{O}}$, since such a policy corresponds to a non-Markovian policy that is approximately optimal for $\mathcal{M}_{f}$. In particular, an approximately optimal policy for $\mathcal{M}_{\text{O}}$ can be seen as a non-Markovian policy for $\mathcal{M}_{f}$ that compresses the history up to any timestep into a running occupancy. This result demonstrates that maintaining the running occupancy up to any timestep is sufficient to achieve optimal behavior in the single-trial truncated-horizon regime equation 8.
@@
Since the class of policies $\Pi_{\textnormal{S}}^{\textnormal{D}}$ suffices for optimality in standard MDPs (Puterman, 2014), we know that at least one stationary deterministic policy $\pi_{\text{O}}\in\Pi_{\textnormal{S}}^{\textnormal{D}}$ for $\mathcal{M}_{\text{O}}$ satisfies $J_{\text{O}}(\pi_{\text{O}})-J_{\text{O}}^{*}=0$, i.e., $\pi_{\text{O}}$ is optimal for $\mathcal{M}_{\text{O}}$. In light of Theo. 2 and Prop. 2, this implies that the corresponding non-Markovian policy $\pi\in\Pi_{\textnormal{NM}}$ for $\mathcal{M}_{f}$, which is deterministic, satisfies $\mathrm{OptGap}_{H}(\pi)=0$, i.e., $\pi$ is optimal for $\mathcal{M}_{f}$. Thus, we can focus our attention on deterministic non-Markovian policies when solving $\mathcal{M}_{f}$ with objective $F_{1,H}$, i.e., for any GUMDP and horizon $H\in\mathbb{N}$, it holds that $\min_{\pi\in\Pi_{\textnormal{NM}}}F_{1,H}(\pi)=\min_{\pi\in\Pi_{\textnormal{NM}}^{\textnormal{D}}}F_{1,H}(\pi).$
@@
3.4 Hardness result for policy optimization in the single-trial regime
@@
(Complete proof in Appendix B.6) We reduce the subset sum problem to the policy existence problem in GUMDPs with objective $F_{1,H}$. The subset sum problem asks whether, given a set $\mathcal{N}=\{n_{0},n_{1},\ldots,n_{N-1}\}$ of $N$ non-negative integers and a target sum $k\in\mathbb{N}$, there exists a subset of $\mathcal{N}$ whose elements sum to $k$. We map every instance of the subset sum problem as a GUMDP such that at each state $s_{i}$, for $i\in\{0,\ldots,N-1\}$, the policy $\pi\in\Pi_{\text{NM}}^{\text{D}}$ needs to decide between selecting: (i) $a_{\text{include}}$, thereby including $n_{i}$ in the sum; or (ii) $a_{\text{not-include}}$, thereby excluding $n_{i}$ from the sum. Then, we set $H\geq N$ and let $f({\bm{d}})=({\bm{n}}^{\top}{\bm{d}}-k)^{2}$, where $\bm{d}$} denotes a discounted occupancy that captures information regarding the actions selected by the agent at each state $s_{i}$. We construct vector ${\bm{n}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ such that ${\bm{n}}^{\top}{\bm{d}}$ equals the sum of the numbers selected by the policy. With this construction, the objective satisfies $f({\bm{d}})=0$ if and only if the sum of the selected numbers equals $k$. By setting $\lambda=0$, we are asking whether there exists a policy such that $F_{1,H}(\pi)\leq 0$. Since $f({\bm{d}})=0$ if and only if the selected numbers sum to $k$, the reduction is complete. ∎
@@
With the above hardness result in mind, we next explore how to develop practical planning algorithms for our problem. Naturally, in a worst-case sense, these algorithms may require a non-polynomial number of steps to retrieve the optimal policy. Nevertheless, our results show it is possible to develop practical algorithms that are superior in comparison to relevant baselines.
@@ 4 Online Planning for GUMDPs in the Single-Trial Regime
As previously shown, solving a GUMDP in the single-trial setting is closely related to solving a corresponding occupancy MDP. This connection allows us to employ an online planning approach in which, at any timestep $t\in\{0,\ldots,H-1\}$ of the interaction with the occupancy MDP, the algorithm receives the current state $\{s_{t},{\bm{o}}_{t}\}$. The online planner then expands a look-ahead search tree where the root node corresponds to state $\{s_{t},{\bm{o}}_{t}\}$. After a given number of iterations, the planning algorithm selects an action to execute in the environment; depending on the selected action, the environment evolves to a new state, and the process repeats until timestep $H$. This online planning strategy is particularly effective, as it allows computational resources to be focused on computing (approximately) optimal actions only along the specific trajectory experienced by the agent. This avoids the prohibitive cost of computing an optimal policy for every state of the occupancy MDP.